OpenStudy (anonymous):

how do i simplify the square root of 12?

jimthompson5910 (jim_thompson5910):

which perfect squares are factors of 12?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so 12 = 4*3

jimthompson5910 (jim_thompson5910):

which means \[\Large \sqrt{12} = \sqrt{4*3}\] \[\Large \sqrt{12} = \sqrt{4}*\sqrt{3}\] see how to finish up?

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

the square root of 4 is not 4

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

good, \[\Large \sqrt{12} = 2\sqrt{3}\]